PA Mo-0StNm MO L-15m E- 2.10 MP P=02 Suci 2) FR distablaltdas)) 3 6 HM DeY May

force reactions
Answer #1

As asked by you, I have calculated force reactions only.

Please rate the answer. Thank you!

Free body diagram: M=0.5 kN. Y) B A RB RA 1.5m EMA=O 0.5+ (-RB) X1.5=0 0.5 1 R8- 1.5 06667 kN 0.5+ CRA) X 1.5=0 -0.6667 kN 0.